In a random sample of 500 observations, 290 successes and 210 failures were detected.

1. Calculate the point estimate of the population proportion of success.

2. Estimate with 95% confidence the population proportion of success.

Use a hypothesis test with significance level 0.05 to test whether the population proportion of success is less than 0.6

● state the null and alternative hypothesis
●state and calculate the appropriate test statistic
●determine the critical value of the test and state the test and state rejection region
●draw an appropriate conclusion
●state whether or not you reject the null hypothesis
1
Expert's answer
2020-03-31T11:17:10-0400

1) Point estimate of population proportion of successes "\hat{p}=\frac{290}{500}=0.58."

2) "n\hat{p}=500\cdot 0.58\geq10,""n(1-\hat{p})=500(1-0.58)\geq10". So we can approximate the binomial distribution with a normal distribution.

"z-value=1.96\\
\hat{q}=1-\hat{p}=1-0.58=0.42.\\
EBP=z-value\sqrt{\frac{\hat{p}\hat{q}}{n}}=1.96\sqrt{\frac{(0.58)(0.42)}{500}}\approx 0.043.\\
(\hat{p}-EBP;\hat{p}+EBP)\\
(0.58-0.043;0.58+0.043)\\
(0.537;0.623)\text{ --- our confidence interval}."

3) "H_0: p=0.6, H_1:p<0.6" (left-tailed test)

"pn=(0.6)500>5\\
(1-p)n=(0.4)500>5"

So we will use z-test.

"z-stat=\frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}}=\frac{0.58-0.6}{\sqrt{\frac{(0.6)(0.4)}{500}}}\approx -0.9129.\\
\Phi(z_{cr})=\frac{1-2\alpha}{2}=0.45.\\
z_{cr}=1.64.\\
(-\infty,-1.64)\text{is the critical region}.\\
z-stat\text{ is not in the critical region. So we accept, do not reject } H_0."